OpenStudy (anonymous):

medal&fan again asap :) The legs of a right triangle are 3 units and 8 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth. // 8.54 units 9.54 units 11.00 units 24.00 units

OpenStudy (godlovesme):

r u familiar with the eqn \[a^2+b^2=c^2\]

OpenStudy (godlovesme):

@KraziiKai r u there?

OpenStudy (godlovesme):

\[3^2+8^=c^2\] \[9+64=c^2\]

OpenStudy (godlovesme):

√73=√c c=8.54 hope that helped :)

OpenStudy (anonymous):

I'm here now, & woah thats confusing :p @Godlovesme

OpenStudy (godlovesme):

lol :') r u familiar with the eqn i posted above^^^^

OpenStudy (anonymous):

@Godlovesme i just watched a video about this, i understand now, thank you for your help!

OpenStudy (godlovesme):

u r welcome glad that helped thanks for the medal :)